OpenStudy (anonymous):

2c2 + 4cd + 3c + 3d i need to facter this through grouping can i and how?

OpenStudy (anonymous):

3d or 6d? if 6d then: 2c^2 + 4cd + 3c + 6d = (2c^2 + 4cd) + (3c + 6d) = 2c(c+2d) + 3(c+2d) =(c+2d)(2c+3)

OpenStudy (anonymous):

OHHH OK I GET THANKS!